# S2C2A: A Flexible Task Space Planning and Control Strategy for Modular Soft Robot Arms

### v1

@@ Abstract
Modular soft robot arms (MSRAs) are composed of multiple independent modules connected in a sequence. Due to their modular structure and high degrees of freedom (DOFs), these modules can simultaneously bend at different angles in various directions, enabling complex deformation. This capability allows MSRAs to perform more intricate tasks than single-module robots. However, the modular structure also induces challenges in accurate planning, modeling, and control. Nonlinearity, hysteresis, and gravity complicate the physical model, while the modular structure and increased DOFs further lead to accumulative errors along the sequence. To address these challenges, we propose a flexible task space planning and control strategy for MSRAs, named $S2C2A$ (State to Configuration to Action). Our approach formulates an optimization problem, $S2C$ (State to Configuration planning), which integrates various loss functions and a forward MSRA model to generate configuration trajectories based on target MSRA states. Given the model complexity, we leverage a biLSTM network as the forward model. Subsequently, a configuration controller $C2A$ (Configuration to Action control) is implemented to follow the planned configuration trajectories, leveraging only inaccurate internal sensing feedback. Both a biLSTM network and a physical model are utilized for configuration control. We validated our strategy using a cable-driven MSRA, demonstrating its ability to perform diverse offline tasks such as position control, orientation control, and obstacle avoidance. Furthermore, our strategy endows MSRA with online interaction capability with targets and obstacles. Future work will focus on addressing MSRA challenges, such as developing more accurate physical models and reducing configuration estimation errors along the module sequence.
@@ I Introduction
In comparison to rigid robots, soft robots have a number of advantages, with their most notable being their flexibility and compliance, which can provide safe yet robust environmental interactions. Due to this, they have been leveraged for various applications, including medical scenarios[1, 2, 3] and bionic robots[4, 5, 6]. Most soft robots are constructed from elastomeric materials[7, 8], incorporate deformable structures[9] or origami-inspired structure[10]. These materials and designs contribute to the lower stiffness and enhanced safety of soft robots compared to their traditional rigid counterparts. Resulting from their compliance, soft robots exhibit high degrees of freedom (DOFs) and, hence, are able to perform more diverse environmental interactions[11]. Consequently, soft robots have been successfully deployed in constrained working environments[12], such as the abdomen[13] and cluttered pipe environments[14].
@@
Among the various categories of soft robots, modular soft robot arms (MSRA) have demonstrated significant potential, particularly due to their higher DOFs than single-module robots. Typically, one MSRA is composed of multiple independent modules connected in series, for example, as utilized by the robots described in [15, 16, 9], where modules are replicated to create a longer, more capable system [17, 16]. In most works, this kind of arm will be named as ’soft continuum robot’[18] or ’soft continuum arm’[19], but we use ’modular soft robot arm’ to emphasize its modularity, which is the focus of this work. The presence of multiple modules increases the degrees of actuation within the same robot length, providing the robots with more diverse shape patterns[9] and a richer working space[16]. Due to these unique advantages, MSRAs have been leveraged in challenging tasks such as obstacle avoidance[20], rotating a handwheel, and opening a drawer[21].
Along with the high DOFs of MSRAs, their complex structures also pose challenges in accurate planning and control due to their softness and modulraity. The inherent nonlinearity and hysteresis of the underlying structures make robot control a complex problem [22]. Furthermore, the modular structure complicates the mapping relationships among the task, configuration, and actuation spaces. As illustrated in Figure1-(A), the module configurations will affect each other along the module sequence and ultimately determine the overall state of the soft robot, which infers that all configurations and their mutual influence should be considered in MSRA control. Such a requirement complicates the MSRA control issue compared to the single module arm control. Although physical model approaches have been widely employed in single module arm problems, the module improvement dramatically complicates the robot models[23], furthermore challenging the proposal of model-based controllers, specifically real-time ones.
@@
Neural networks (NNs), particularly recurrent neural networks (RNNs), have been extensively leveraged in soft robot modeling and control. Due to the nonlinear activation functions, NNs are suitable for handling the nonlinear motions characteristic of soft robots[15]. Moreover, the recurrent structures of RNN can be employed to address sequence-related issues such as time delay[24] and MSRA configuration control[16]. As a result, RNNs have the potential to tackle the complexities posed by MSRAs.
@@
This paper aims to propose a flexible MSRA task space planning and control strategy named S2C2A, as illustrated in Figure 1-(C). A configuration planning approach S2C is employed to generate configuration trajectories based on the task requirements online or offline using an optimization problem, and we apply a biLSTM network as the MSRA forward model in optimization. Additionally, a feedback configuration controller C2A specifically designed for MSRAs is employed to execute configuration control tasks, and both a biLSTM network and a physical controller are utilized. We validate our approach on a cable-driven MSRA, demonstrating its flexibility in achieving a variety of tasks.
@@
    We introduce a hierarchical MSRA task space planning and control strategy named S2C2A. The configuration trajectory is planned at first, and a configuration controller is used to control the configuration.
@@
    We leverage both biLSTM and physical controller for configuration control even in the presence of inaccurate internal sensing feedback instead of external sensing.
  4. 4.
@@
    Real experiments have been carried out to validate our strategy. Offline Tasks including position control, orientation control, and obstacle avoidance have been achieved successfully. Moreover, online interactions with the target and obstacle are performed.
@@
Figure 1: (A) Modular soft robot arm motion diagram. The action of each module $A_{*}$ (green) will directly affect the configurations $C_{*}$ of each module (red) and indirectly affect the whole robot system state $S_{*}$ (blue). The configurations of modules are affected by gravity and the adjacent modules. (B) MSRA applied in our experiments can deform in an S-shape, which single-module robots cannot achieve. (C) S2C2A comprises a configuration planning strategy S2C and a configuration controller C2A. S2C can propose the target configuration online or offline based on the task request. C2A collects configuration estimation from internal sensors and determines action based on the sensor feedback and target configuration from S2C.
@@
The remainder of the paper is structured as follows: Section II reviews works related to soft robot planning and control approaches, with a focus on MSRA. Section III describes the robot setup and the definitions of configuration and state. Section IV details the task space planning and control strategy S2C2A, covering configuration planning method S2C using biLSTM and configuration controller based on either bilSTM or a curvature model. Section V presents the real experimental results, which demonstrate that our controllers can perform a range of tasks, from basic ones like position and orientation control to more complex ones like obstacle avoidance and online target following. Section VI summarizes this paper and discusses potential directions for future work.
II Related Work
II-A MSRA Planning
In robot planning, trajectories in the task or configuration space are calculated to achieve task requirements like trajectory following, orientation control, and obstacle avoidance. Effective planning is essential for high DOF robots like MSRA[25] and challenging working spaces like obstacle-included[26] and constrained environments[12]. Most studies about soft robot planning rely on physical models, and the constant curvature (CC) model[12] is the most popular one due to its simplicity and low computational cost. The shapes of the robot backbone, inflated chambers, and obstacles are considered in the planning strategy of [12] to propose available robot shapes for the MSRA. Similarly, the curvature model is also utilized in [27] in conjunction with rapidly exploring random trees for planning. To achieve inspection tasks in cluttered space, one MSRA is segmented into independent curvatures step by step to find feasible shapes in [14]. Additionally, a MSRA is represented by Bézier curves in [28] for planning.
In addition to physical models, some data-driven approaches are utilized. For instance, the Gaussian Mixture Model (GMM) is deployed in [25] to encode the spatial and temporal values from the human demonstration, and trajectory is generated through regression for MSRA. A 3D neural network is leveraged in [29] to reduce unnecessary sweeping motion in robot-assisted surgeries. The improved artificial potential field method is leveraged in [30] for an MSRA with five modules, where constraints and targets are integrated into loss functions of planning optimization. However, this optimization-based approach is only validated in simulation and with a single type of task. Similarly, optimization techniques are also employed in [31, 32], where the distances between the current state and the targets are defined as rewards, and those between the current state and the obstacles are defined as penalties.
It is challenging to build an accurate physical model for MSRAs, while data-driven approaches like NNs have limitations regarding task adaptability. Generally, a single NN is designed to handle a specific task, and it is difficult to adapt to varying or additional requirements with the same trained model. For instance, the NN controller in [33] aims to control robot position and orientation simultaneously. Utilizing such a controller, it is not possible to set only the position target without putting constraints on orientation, nor can the trained controller be easily extended to additional functions such as comprehensive shape control. Although optimization and NN forward model are also used in [24] for MSRA, it only achieves one kind of 2D trajectory following tasks and does not take full advantage of this optimization planning strategy. In this case, we leverage a biLSTM as a forward model within an optimization framework to enable flexible configuration planning. By modifying the cost function in the optimization problem, configuration trajectories tailored to different tasks can be determined.
II-B Soft Robot Control
Based on proposed trajectories in the task or configuration space, a proper controller is essential to achieve trajectory following tasks. There have been several reviews focusing on soft robot control [22, 34, 35, 36]. Jacobian controllers have been applied in soft robots[37]. The Jacobian matrix is updated in real time, and the action is decided by either optimization problems [37] or the inverse Jacobian matrix[38]. Moreover, various categories of physical models, such as Piecewise Constant Curvature (PCC)[39, 40, 41], Cosserat rod[5, 42, 43], and Finite Element Methods (FEM)[44, 45, 46], have been leveraged for soft robot control. These methods aim to describe robot motion through physical deformation and design controllers based on these physical models.
In addition to the above physical controller, data-driven methods have also been applied to soft robot control. Statistical approaches such as support vector regressor (SVR)[47], Gaussian mixture model (GMM)[48], and Gaussian process regression (GPR)[49, 3] have been utilized. The soft robot motion is summarized by optimizing the parameters to propose the controllers. Currently, NNs are the most popular choice in this area. Considering the nonlinear activation functions, NN controllers are suitable for soft robots exhibiting nonlinear behaviors. The research in soft robot NN control starts with simple NNs like extreme learning machine (ELM)[50, 51] and radial basis function (RBF) [52, 53, 54]. Multilayer perceptron (MLP)[55, 56] has been widely applied in soft robots, including sensing, modeling, and control. Due to the recurrent structure and soft robot hysteresis, RNNs are particularly well-suited for soft robot control[57, 58].
Although NN controllers have been extensively leveraged in many soft robot applications, they may not always address the specific physical requirements of these robots. For instance, the cable displacements in the cable-driven robots are not fully independent[59], and such coupling cannot be represented by the existing NN layers. Additionally, most works utilizing NN in MSRA employ external sensing systems like optical cameras, which are challenging to implement in real applications. In this work, we introduce an additional layer at the end of the biLSTM network to address the physical actuation requirements of the robot. In addition, we incorporate a CC controller to prove the flexibility of our proposed strategy. Configuration feedback is provided by only internal sensors, and the costly external sensing system is not necessary during control.
III Setup and Specification
@@
Figure 2: (A) Robot setup. The cable-driven robot is composed of three independent segments, while each segment is actuated by three motors. The encoder in each motor provides feedback for configuration estimation. The optical tracking system is applied to provide ground truth robot states. (B) Configuration and state. Configurations (red) refer to segment end orientation relative to the segment base, and states (blue) refer to the position and orientation of every segment relative to the robot base. (C) Module actuation. The module configuration can be calculated by module bending angle $\phi$ and neutral surface direction $\theta$, which can be estimated by the displacements of cables $c_{1},c_{2},c_{3}$.
III-A Experimental Setup
In this paper, we utilized a cable-driven modular soft robot arm for our experiments, as shown in Figure 2-(A). The robot comprises three independent modules (about 0.2m for each module), and each module is actuated by three motors (DYNAMIXEL xl330-m077-t) fixed on the MSRA base via three cables placed at an interval angle of 120 along the circumferential direction. The cables of modules II and III pass through housing tubes, guaranteeing each module’s independence. In other words, the actuation of cable III will only affect module III instead of the whole MSRA due to the housing tubes. More details of the robot setup, such as MSRA material and mechanical properties, can be found in [9].
@@ III-A Experimental Setup
For capturing a ground truth for the robot’s position, several optical track markers are fixed at the end of each module and robot base. Six optical tracking cameras (Optitrack Prime 13) and motion capture software Motive (Optitrack) collect each module end positions and orientations for NN training and only monitoring during control. A laptop (Ubuntu 20.04, CPU i5-12500H, and GPU RTX 3050) receives optical tracking information from the Motive computer via an ethernet switch and communicates with motors via one u2d2 power hub board (DYNAMIXEL). All devices run at about 10Hz in data collection and control tasks.
III-B State and Configuration
@@
As shown in Figure 2-(B), we utilized the unit end orientation vector of each module subject to the module end $(o_{x},o_{y},o_{z})$ to represent module configuration $C$. We utilize it to describe module configuration following [16] rather than bending directions and angles because such a representation is more suitable for NN data training as the vector length has a fixed range and changes continuously. In contrast, bending direction angles experience abrupt changes from 360 to 0 when the orientation completes a full rotation, making them unsuitable for NN data training. The robot state $S$ is composed of the position $(p^{\prime}_{x},p^{\prime}_{y},p^{\prime}_{z})$ and orientation $(o^{\prime}_{x},o^{\prime}_{y},o^{\prime}_{z})$ of every module end subject to the robot end. Therefore, module configurations are independent, and the configuration of one module will not affect those of the other modules. In contrast, the state of one module is affected by the states of base modules and its own configuration.
III-C Configuration Estimation

### v2

# A Versatile Neural Network Configuration Space Planning and Control Strategy for Modular Soft Robot Arms

@@ Abstract
Modular soft robot arms (MSRAs) are composed of multiple modules connected in a sequence, and they can bend at different angles in various directions. This capability allows MSRAs to perform more intricate tasks than single-module robots. However, the modular structure also induces challenges in accurate planning and control. Nonlinearity and hysteresis complicate the physical model, while the modular structure and increased DOFs further lead to cumulative errors along the sequence. To address these challenges, we propose a versatile configuration space planning and control strategy for MSRAs, named $S2C2A$ (State to Configuration to Action). Our approach formulates an optimization problem, $S2C$ (State to Configuration planning), which integrates various loss functions and a forward model based on biLSTM to generate configuration trajectories based on target states. A configuration controller $C2A$ (Configuration to Action control) based on biLSTM is implemented to follow the planned configuration trajectories, leveraging only inaccurate internal sensing feedback. We validate our strategy using a cable-driven MSRA, demonstrating its ability to perform diverse offline tasks such as position and orientation control and obstacle avoidance. Furthermore, our strategy endows MSRA with online interaction capability with targets and obstacles. Future work focuses on addressing MSRA challenges, such as more accurate physical models.
@@ I Introduction
Most soft robots are made of elastomeric materials [1] or built using deformable structures [2]. These materials and structures contribute to the lower stiffness and enhanced safety of soft robots compared to their traditional rigid counterparts. Due to their safety, they have been leveraged for various applications, including medical scenarios [3]. Also, resulting from their compliance, soft robots exhibit high degrees of freedom (DOFs) and, hence, are able to perform environment exploration tasks[4]. Consequently, soft robots have been successfully deployed in complex working environments, such as the abdomen [1] and cluttered pipe environments [5].
@@
Among the various categories of soft robots, modular soft robot arms (MSRA) have demonstrated significant potential, particularly due to their higher DOFs than single-module robots. Typically, one MSRA is composed of multiple independent modules connected in series, which can be different [6] or uniform [7]. “Soft continuum manipulators’ [4] or ’continuum robots’ [8] can contain multiple modules and also fall into the MSRA category. We use ’modular soft robot arm’ instead of these general and common names to emphasize modularity, which is the focus of this work. The presence of multiple modules increases the degrees of actuation within the same robot length, providing the robots with richer working spaces and more diverse shape patterns [2]. Due to these unique advantages, MSRAs have been leveraged in challenging tasks such as obstacle avoidance [8] and challenging environment exploration [9].
Along with the high DOFs, MSRAs also possess some difficulties unique to single-module soft robots, mainly induced by the complicated mappings among the task, configuration, and actuation spaces. In single-module robot research, most works map directly between the task and actuation space, and the configuration space is ignored [10]. However, the module configurations play a significant role in constructing the robot shape in the task space, as shown in Fig. 1(A). Various module features can be applied as the configuration, like bending angle and direction [5] and module end orientation [7]. To measure the MSRA configuration, external sensors like the optical tracking systems [11] and internal sensors like inertial measurement unit (IMU) [12] and motor encoder [2] are leveraged. However, most internal sensing approaches suffer from low accuracy caused by cable slack and drift [12]. Targeted and intelligent algorithms are required to cope with sensing inaccuracy.
Furthermore, the mapping between the task and configuration space induces redundancy in MSRA control. Most control targets, like position and orientation, can be managed by multiple configuration choices [11]. This kind of redundancy endows MSRA with adaptability to various high-level tasks, such as obstacle avoidance. Meanwhile, proper algorithms are excepted to determine the configuration choice and fully achieve this potential. Additionally, modularity also induces difficulty in configuration control. In single-module robot research, although delay, hysteresis, and nonlinearity have already posed challenges in accurate control, the robot state basically relies on the previous states and actions. In addition to these challenges, configuration control in MSRA should also consider the influence of the neighboring modules [7] as shown in Fig. 1(A). Overall, MSRA control requests an unique configuration planning strategy and a configuration controller.
TABLE I: Related works of MSRA
Ref. Task Module number Dimension Model Planning Details
[4] position 6 2 PCC forward model + optimization confined environment exploration
[13] position 4 2 pseudo rigid robot model forward model + optimization target reaching
[14] position 2 3 PCC artificial potential field obstacle avoidance
[5] position 3 3 PCC inverse model + sampling simulation, cluttered space exploration
[9] position 6 2 PCC GMR learning from demonstration, passing holes
[15] position 3 3 PCC RRT obstacle avoidance
[16] position + orientation 2 3 PCC forward model + optimization obstacle avoidance
[2] position + orientation 3 3 PCC forward model + optimization interaction, our previous work
[8] position + orientation 3 3 PCC multiple configuration solutions + optimization obstacle avoidance
[11] position + orientation 3 3 PCC fuzzy logic controller close loop for disturbance rejection
[17] shape 3 3 PCC user-defined curve + optimization fitting drawn curve
[18] shape 2 3 Biarc and Bézier Curve inverse model passing tube
Ours shape 3 3 NN NN + optimization internal sensing
Works related to MSRA planning and control are summarized in Table I. Similar to single-module robot control, the end position is the most common control target. To cope with the challenging mappings among the task, configuration, and actuation spaces, optimization based on a forward model is an effective solution. For instance, a planar MSRA is modeled as Piecewise Constant Curvature (PCC) in [4], and one optimization approach is employed to plan proper configurations based on task requirements like position control and collision avoidance. Similarly, optimization in [13] directly maps from the task space to the actuation space based on the pseudo-rigid robot model. In addition to the optimization, sampling based on the inverse model can also contribute to configuration planning, and this strategy can achieve tasks like cluttered space exploration [5]. Rapidly exploring random trees (RRT) samples the configuration space in [15] and proposes feasible trajectories. The Gaussian Mixture Model (GMM) is deployed in [9] as the inverse model from task space to actuation space and can generate trajectory learning from the demonstration.
In addition to the position control task, some unique tasks can only be achieved by MSRAs, such as position and orientation control [2] and shape control [17]. Similar to the position control strategy, forward models like PCC [16] and optimization are utilized for configuration planning based on target position and orientation. Inverse kinematics is also applied to propose multiple configuration solutions according to the target position and orientation [8]. Additionally, the fuzzy logic controller [11] directly maps between the task and actuation space and can keep end position/orientation invariant while changing orientation/position. Beyond end position and orientation control, the MSRA shape can also be controlled, fully achieving the potential of its structure. Optimization [17] and inverse model [18] are also leveraged for shape control.
@@
Considering the inaccurate internal configuration estimation and the challenges of MSRA control, we leverage neural networks (NNs) in this work. NNs, particularly recurrent neural networks (RNNs), have been extensively leveraged in soft robot modeling and control. Due to the nonlinear activation functions, NNs are suitable for handling the nonlinear motions characteristic of soft robots. Moreover, the recurrent structures of RNN can be employed to address sequence-related issues such as time delay [19] and MSRA module sequence [7]. Several reviews have also highlighted the effectiveness of this data-driven approach in soft robot modeling and control [20, 21]. In addition, as a data-driven approach, RNN has the ability to map between spaces without explicit physical relationships, like the MSRA task space and the configuration space constructed by the inaccurate sensing feedback, as shown in Fig. 1(C). Therefore, we believe that RNNs have the potential to tackle the complexities posed by MSRAs.
@@
This paper aims to propose a versatile MSRA configuration space planning and control strategy named S2C2A based on NNs and internal sensing feedback, as illustrated in Fig. 1(C). Internal sensors provide rough sensing feedback to represent module configuration. A configuration planning approach S2C via optimization is employed to generate configuration trajectories based on the task requirements online or offline, and we apply a biLSTM network as the MSRA forward model in optimization. Additionally, a biLSTM configuration controller C2A is designed specifically for MSRAs, and the rough internal configuration feedback is employed to perform configuration control tasks. We validate our approach on the cable-driven MSRA shown in Fig. 1(B) and compare it with our previous work [2], demonstrating its versatility by achieving a variety of tasks.
@@
    We leverage a biLSTM controller for configuration control to deal with the inaccurate configuration estimation from the internal sensing.
@@
    We carried out real experiments to validate the accuracy and versatility of our strategy. Offline tasks, including position control, orientation control, and obstacle avoidance, have been achieved successfully. Moreover, online interactions with the target and obstacle are performed.
@@
Figure 1: (A) Modular soft robot arm motion diagram. The action of each module $A_{*}$ (green) will directly affect the configurations $C_{*}$ of each module (red) and indirectly affect the whole robot system state $S_{*}$ (blue). The configurations of modules are also affected by gravity and the adjacent modules. (B) MSRA applied in our experiments can deform in an S-shape, which cannot be achieved by single-module robots. (C) S2C2A comprises a configuration planning strategy S2C and a configuration controller C2A. NN can contribute to the S2C strategy and, based on the various task requests, propose the target configuration online or offline. The other NN can serve as a C2A controller, collecting configuration estimation from internal sensors and determining action based on sensor feedback and target configuration from S2C. This approach can achieve accurate MSRA state control without the knowledge of real configurations.
@@
The remainder of the paper is structured as follows: Section II describes the robot setup, the definitions of configuration and state, the configuration estimation approach, and data collected from our MSRA. Section III details the task space planning and control strategy S2C2A, covering configuration planning method S2C using biLSTM and configuration controller based on inaccurate internal sensing feedback. Section IV presents the real experimental results, which demonstrate that our controllers outperform our previous approach [2] and can perform a wide range of tasks, from basic ones like position and orientation control to more complex ones like obstacle avoidance and online target following. Section V summarizes this paper and discusses potential directions for future work.
II Setup and Specification
@@
Figure 2: (A) Robot setup. The cable-driven robot is composed of three independent segments, while each segment is actuated by three motors. The encoder in each motor provides feedback for configuration estimation. The optical tracking system is applied to provide ground truth robot states and configurations. (B) Configuration and state. Configurations (red) refer to segment end orientation relative to the segment base, and states (blue) refer to the position and orientation of every segment relative to the robot base. (C) Module actuation. The module configuration can be calculated by module bending angle $\phi$ and neutral surface direction $\theta$, which can be estimated by the displacements of cables $c_{1},c_{2},c_{3}$.
II-A Experimental Setup
In this paper, we utilized a cable-driven modular soft robot arm for our experiments, as depicted in Fig. 2(A). The robot comprises three independent modules (approximately 0.2 meters per module) based on Trimmed Helicoid (TH) structures. Each module is actuated by three motors (DYNAMIXEL xl330-m288-t) mounted on the MSRA base and connected via three cables arranged at 120° intervals along the circumferential direction. In this paper, we apply $I,II,III$ to represent the module number and $1\sim 5$ to represent the time step. The cables of Modules I and III are positioned at 0°, 120°, and 240°, while the cables of Module II are positioned at 60°, 180°, and 300°. The cables of Module I are directly attached to the motors with 20mm diameter pulley wheels. The cables of Modules II and III traverse housing tubes ($4\times 2$mm PTFE tubes), thereby guaranteeing each module’s independence.
The housing tube-cable structure in our robot shares the same working principle as the bicycle brake system. The cable inside can transmit the displacement between two ends while not deforming the outer tube. Similarly, the housing tubes in our MSRA are fixed at both ends—extending from the base of the motor to the base of each module—while passing through pre-designed channels embedded within the TH structure. The force exerted by the internal tendons on the housing tubes is counterbalanced at both ends by the compressive force at the corresponding sections. As a result, the housing tubes effectively transmit the tension force directly from the motor base to a specific soft robotic section. Consequently, the tendon force influences only the targeted section without affecting other sections. The only external factors influencing adjacent sections are the friction between the outer surface of the housing tube and the internal channels of the robot arm, as well as the reactive bending moment induced by arm flexion. However, these effects are negligible in our design due to the use of smooth and highly flexible PTFE tubes as housing tubes. Further details regarding the robot setup, including the MSRA material and mechanical properties, can be found in [2] and the Supplementary Video.
@@ II-A Experimental Setup
For capturing a ground truth for the robot’s position, several optical track markers are fixed at the end of each module and robot base. Six optical tracking cameras (Optitrack Prime 13) and motion capture software Motive (Optitrack) collect each module end positions and orientations for NN training and only monitoring during control. A laptop (Ubuntu 20.04, CPU i5-12500H, and GPU RTX 3050) receives optical tracking information from the Motive computer via an ethernet switch and communicates with motors via one u2d2 power hub board (DYNAMIXEL). The whole system runs at about 10Hz in data collection and control tasks.
II-B State and Configuration
@@
As shown in Fig. 2(B), we utilized the unit end orientation vector of each module subject to the module end $[o_{x}\ o_{y}\ o_{z}]\in R^{3}$ to represent module configuration $C$. We utilize it to describe module configuration following [7] rather than bending directions and angles because such a representation is more suitable for NN data training as each element falls into the range of [-1, 1] and changes continuously during the MSRA motion. In contrast, bending direction angles undergo sudden shifts from 360 to 0 upon completing a full rotation, making them unsuitable for NN data training.
The robot state $S=[S_{I}\ S_{II}\ \dots S_{m}]\in R^{6\times m}$ is composed of the states of all modules as shown in Fig. 2(B), where $m$ represents module number and is 3 in this work. For the module $n$, its state $S_{n}=[p^{\prime}_{n}\ o^{\prime}_{n}]^{T}\in R^{6}$ can be represent by its end position $p^{\prime}_{n}=[p^{\prime}_{n,x}\ p^{\prime}_{n,y}\ p^{\prime}_{n,z}]\in R^{3}$ and orientation $o^{\prime}_{n}=[o^{\prime}_{n,x}\ o^{\prime}_{n,y}\ o^{\prime}_{n,z}]\in R^{3}$ subject to the MSRA base. Overall, module configurations $C$ are independent, and from the geometry perspective, the configuration of one module does not affect the configuration of the other modules. In contrast, the state of one module is affected by the configurations of base modules and its own configuration. For instance, the state of the module $n$, $S_{n}$, is determined by the configurations of the modules $1\sim(n-1)$ and its own configuration.
II-C Configuration Estimation